OpenStudy (anonymous):

What is the value of x? 8x-9=3x+16 helppp

OpenStudy (campbell_st):

start by subtracting 3x from both sides of the equation.... when you have done that, add 9 to both sides of the equation. then you should be able to solve for x

OpenStudy (anonymous):

idk how to do that @campbell_st

OpenStudy (dmndlife24):

8x-9=3x+16 (+9) (+9) 8x=3x+25 5x=25 x=5

OpenStudy (campbell_st):

8x - 9 = 3x + 16 -3x -3x ----------------- 5x - 9 = 16 now add 9 to both side to handle the -9

OpenStudy (anonymous):

is it x=5 @campbell_st